# Pasadena, CA 91125, USA

### v1

@@ 3 Methods
    Extract the 3d $\mathcal{N}=2$ Abelian Chern-Simons matter theory from the above expression for the half-index.
@@
    This gauge theory has a superpotential, which triggers an RG flow to $\mathcal{N}=4$ SCFT in the IR. The superpotential breaks most of the symmetry except for the distinguished $U(1)_{A}$ symmetry (and flavor symmetries of 4d), which becomes part of the $SO(4)_{R}$ symmetry. Find this $U(1)_{A}$ symmetry.
@@
    Refine the half-index by turning on fugacity $T$ for $U(1)_{A}$. This gives the Macdonald index of the 4d theory.
@@
The starting point of our procedure is an observation that the fermionic sum formula from the IR formula can be viewed as a twisted half index of a 3d $\mathcal{N}=2$ Abelian Chern-Simons matter theory (ACSM) Gaiotto:2024ioj . The half-index of a 3d $\mathcal{N}=2$ theory with a supersymmetric boundary condition $B$ is defined by Gadde:2013wq ; Gadde:2013sca ; Yoshida:2014ssa ; Dimofte:2017
@@
Consider a $U(1)^{N}$ Chern-Simons theory with level matrix $K$ and $M$ chirals $\Phi_{\alpha}$ with $U(1)_{i}$ gauge charge $Q_{i\alpha}$, and $R$-charge $R_{\alpha}$. Considering the supersymmetric Dirichlet boundary conditions to all the vector and chiral multiplets, the half-index reads Dimofte:2017
$I\!\!I\!(q,{\boldsymbol{x}})=\frac{1}{(q;q)_{\infty}^{N}}\sum_{\mathbf{m}\in\mathbb{Z}^{N}}q^{\frac{1}{2}\mathbf{m}^{t}K\mathbf{m}+\frac{1}{2}\boldsymbol{\rho}\cdot\mathbf{m}}{\boldsymbol{x}}^{K{\bf m}}\prod_{\alpha=1}^{M}\left(q^{R_{\alpha}/2}q^{1-{\mathbf{m}}\cdot Q_{\alpha}}{\boldsymbol{x}}^{-Q_{\alpha}};q\right)_{\infty}\ ,$ (40)
where $\rho_{i}$ denotes the mixed Chern-Simons level between the gauge group $U(1)_{i}$ and background $U(1)_{R}$ symmetry, and $x_{i}$’s are fugacity for the boundary current $U(1)^{N}_{\partial}$. In most of the examples below, we consider the fully specialized index with ${\boldsymbol{x}}\rightarrow 1$.
Upon comparing this expression with the one obtained from the IR trace formula for the Schur index (23), we extract an abelian Chern-Simons theory. For example, the Schur index of the form (27) for the $(A_{1},A_{2N})$ theory gives rise to $U(1)^{2N}$ Chern-Simons (CS) theory with where to each $U(1)$ with coupled with 2 chiral multiplets of charge $1$, so that we have in total $M=4N$ chiral multiplets. We can also read the CS level matrix $K$, mixed U(1)RU(1)-R CS level $\rho$ for each $U(1)$ gauge factors, and $R$-charges for the chiral multiplets by comparing the formulae.
This description of the abelian Chern-Simons theory can be motivated by considering a Janus configuration in the Coulomb branch of the 4d SCFT, induced by the $U(1)_{r}$ twisting Gaiotto:2024ioj .
@@
The abelian Chern-Simons theory obtained in this way is expected to flow to a SCFT with $\mathcal{N}=4$ supersymmetry enhancement, after a suitable superpotential deformation. The IR $\mathcal{N}=4$ SCFT has an $\mathcal{N}=2$ superconformal subalgebra where $U(1)_{R}$ symmetry embeds into the full $\mathcal{N}=4$ R-symmetry group $SU(2)_{H}\times SU(2)_{C}$, leaving $U(1)_{A}=U(1)_{H}-U(1)_{C}$ as a distinguished global symmetry. In the case where the 3d $\mathcal{N}=4$ SCFT is expected to have a zero-dimensional Coulomb/Higgs branch, we anticipate that there exists a suitable superpotential deformation that lifts most of the global symmetry of the UV ACSM, leaving one combination of $U(1)$, which can be identified with the IR $U(1)_{A}$ symmetry. To find the superconformal $R$ symmetry, we consider a one-parameter mixing of the $R$ symmetry
@@
and perform $F$-maximization Jafferis:2010un . We define our normalization so that $\nu=0$ corresponds to the superconformal $R$ symmetry $U(1)_{R_{0}}=U(1)_{R}=U(1)_{C}+U(1)_{H}$.
@@
In the special limit $\nu=-1$, the half-index calculates the index with the twisted spin $J+R_{\nu}/2=J+R_{H}$, realizing the topological $A$-twist. We therefore define
@@
The 3d $\mathcal{N}=4$ SCFT obtained by twisted compactification of $(A_{1},G)$ Argyres-Douglas theories are expected to have zero-dimensional Coulomb branches, while their Higgs branches coincide with those of the original 4d SCFT. On the other hand, the proposed $\mathcal{N}=2$ ACSM descriptions generally exhibit a larger global symmetry than the Higgs branch symmetry. We therefore expect that there exists an appropriate superpotential deformation that breaks the global symmetries to $U(1)_{A}\times T_{H}$, where $T_{H}\subset G_{H}$ denotes the Cartan subgroup of the Higgs branch symmetry.
@@
    The superconformal index contains a term $q^{3/2}(T+T^{-1})$, signaling the existence of the extra supercurrent multiplets in $\mathcal{N}=4$ algebra [^2] This condition is neither necessary nor sufficient, since there may exist, for example, chiral primary operators with $R$-charge 3. In fact, we have examples where supersymmetry enhancement is expected, yet this term is partially or entirely absent., as well as a possible moment map operator for the Higgs branch symmetry appearing at $q^{1/2}$ order (when such a symmetry is present).
@@
    The Coulomb branch Hilbert series is trivial. [^3] For a 3d SCFT, the superconformal index can be written as $\mathcal{I}(\mathfrak{q},\mathfrak{t})=\mathrm{tr}_{\mathbb{S}^{2}}(-1)^{F}\mathfrak{q}^{j_{3}+\frac{1}{2}(R_{H}+R_{c})}\mathfrak{t}^{R_{H}-R_{C}}$ (45) and it gets contribution only from states satisfying $E-R_{H}-R_{C}-j_{3}=0$. We can rewrite the index by $\mathcal{I}(\mathfrak{q},\mathfrak{t})=\mathrm{tr}(-1)^{F}(\mathfrak{q}^{\frac{1}{2}}\mathfrak{t})^{E-R_{c}}(\mathfrak{q}^{\frac{1}{2}}\mathfrak{t}^{-1})^{E-R_{H}}.$ (46) The Higgs and Coulomb branch Hilbert series are obtained by taking the limit Razamat:2014 Coulomb: $\qquad\mathfrak{q}\rightarrow 0,\mathfrak{t}\rightarrow 0,\qquad\mathfrak{q}^{\frac{1}{2}}\mathfrak{t}^{-1}\;\text{fixed}$ (47a) Higgs: $\qquad\mathfrak{q}\rightarrow 0,\mathfrak{t}\rightarrow\infty,\qquad\mathfrak{q}^{\frac{1}{2}}\mathfrak{t}\;\text{fixed}$ (47b) which give extra shortening condition $E=R_{H}$ and $E=R_{C}$ for the trace, respectivley.
@@
    The Higgs branch Hilbert series agrees with the Higgs branch chiral ring of 4d SCFT.
In some low-rank examples, this procedure determines the $U(1)_{A}$ symmetry in the ACSM description, and one can check that the refined half-index (43) reproduces the conjectural Macdonald index in section 2.3.
One subtlety in identifying $U(1)_{A}$ arises from the fact that a global symmetry can be redefined by mixing it with gauge symmetries. The Dirichlet half-index is sensitive to this choice because the boundary gauge symmetry is broken to a global transformation. Therefore, the computation of the refined half-index is not completely canonical and depends on this choice. In the examples below, we show that there exists a representation of the $U(1)_{A}$ symmetry that reproduces the conjectural Macdonald index presented in section 2.3, providing a strong consistency check despite its non-uniqueness.
For higher rank examples, where F-maximization is not practically feasible, we begin with the assumption that the Schur index coincides with the A-twisted half-index of 3d $\mathcal{N}=4$ SCFT, and that the conjectural Macdonald index introduced in section 2.3 computes its refined half-index. This assumption completely determines the data for the superconformal R-symmetry. We then compute the corresponding superconformal index and the Coulomb and Higgs branch Hilbert series, and test them against the conditions described above.
@@ 4 Examples
In this section, we perform our procedure for the $(A_{1},G)$ type Argyres-Douglas theories to recover our conjectural closed-form expression for the Macdonald index.
@@ 4.1 $(A_{1},A_{2n})$ theory
The BPS quiver for the $(A_{1},A_{2n})$ theory is given by the figure 2. Cecotti:2010fi
@@
This describes the canonical chamber with an equal number of sources and sinks. The trace formula (23) gives the Schur index Cordova:2015nma
@@
where $b^{g}_{ij}=2\delta_{ij}-C[g]_{ij}$ and $C[g]$ is the Cartan matrix of $g$ (having only mixed CS level $1$ between the $U(1)_{i}$ and $U(1)_{i+1}$). Here we used the identity $(q)_{k}=(q)_{\infty}/(q^{k+1})_{\infty}$. This expression agrees with the character of the Virasoro minimal model $M(2,2n+3)$ Cecotti:2015lab .
Now, we find that this formula is identical to that of the half index of the A-twisted 3d $U(1)^{2n}$ Chern-Simons theory with level
$K=b^{A_{2n}}=\begin{pmatrix}0&1&0&0&\cdots&0&0\\ 1&0&0&0&\cdots&0&0\\ 0&0&0&1&\cdots&0&0\\ 0&0&1&0&\cdots&0&0\\ &&&\vdots\\ 0&0&0&0&\cdots&0&1\\ 0&0&0&0&\cdots&1&0\end{pmatrix}\\ \$. (49)
Also, we have two chiral multiplets $\Phi_{i}$ and $\widetilde{\Phi}_{i}$ with charge 1-1 for each $U(1)_{i}\subset U(1)^{2n}$, giving a total of $4n$ chiral multiplets. Therefore, we can write the matter content and the charge matrix $Q_{i\alpha}$ as
@@
The half-index formula shows that the (twisted) $R$-charges for the chiral multiplets are all zero and the mixed Chern-Simons level for $U(1)_{R}$ and $U(1)_{i}$ is $\rho_{i}=2$ for all $i=1,2,\cdots,2n$. Before adding the superpotential term, the ACS theory has $U(1)^{4n}$ flavor symmetry.
Now our task is to identify the superpotential deformation that breaks all the symmetries except for one, which we denote as $U(1)_{A}$, under which we reproduce the Macdonald grading, and triggers an RG flow to $\mathcal{N}=4$ SCFT in the IR.
@@ 4.1.1 $A_{2}$
Let us start with the simplest case of $n=1$. The Schur index formula (48) gives
@@
This expression is identical to that of the half-index of the 3d $\mathcal{N}=2$ Abelian Chern-Simons theory with $U(1)^{2}$ gauge group, with the following data:
@@
Before turning on superpotential, this gauge theory has $U(1)^{4}$ flavor symmetry, corresponding to the rotations of each chiral multiplet. We have to break $U(1)^{4}$ flavor symmetry down to the single $U(1)_{A}$ by turning on the appropriate superpotential formed out of dressed monopole operators. Such monopole operators should have $R$-charge 2, and there are 4 of them in total, as listed in Table 1.
@@
Consider the following superpotential [^4] We choose arbitrary 3 monopoles out of 4. Any choice would be equivalent thanks to symmetry.
@@
The resulting unbroken global symmetry is $U(1)_{\widetilde{\Phi}_{1}}$, which rotates $\widetilde{\Phi}_{1}$ by +1+1 and keeps the rest intact. By mixing with gauge symmetry, this global symmetry can also be represented as a combination of the topological symmetry for the second gauge node and the symmetry that rotates $\Phi_{1}$ with charge 1-1.  [^5] Consider the bosonic action $2A_{1}dA_{2}+|D_{(-A_{1}+A^{\text{bg}})}\widetilde{\Phi}|^{2}+|D_{(-A_{1})}\Phi|^{2}+\cdots$. By redefining the gauge connection $A_{1}\rightarrow A_{1}+A^{\text{bg}}$, we obtain $2A_{1}dA_{2}+2A^{(\text{bg})}dA_{2}+|D_{(-A_{1})}\widetilde{\Phi}|^{2}+|D_{(-A_{1}-A^{(bg)})}\Phi|^{2}+\cdots$ We will now show that this is the desired superpotential and verify that this unbroken symmetry is the correct axial symmetry $U(1)_{A}$ that accounts for the Macdonald grading.
We can compute the superconformal index using the integral formula with our gauge theory data, refined to include the candidate $U(1)_{A}$ fugacity. If it is indeed the correct $U(1)_{A}$ part of $SO(4)_{R}$ symmetry, this gives the $A$-twisted superconformal index $\mathcal{I}_{SCI}(\mathfrak{q},T)|_{T\rightarrow\mathfrak{q}^{\frac{1}{2}}T}$ since $R$ symmetry is shifted by $U(1)_{A}$. To recover the superconformal index before twisting, we take $T\rightarrow\mathfrak{q}^{-\frac{1}{2}}T$ of the twisted index to obtain
@@
The term $(T+T^{-1})\mathfrak{q}^{\frac{3}{2}}$ (as an $\mathcal{N}=2$ index) arises from an extra supercurrent multiplet giving rise to $\mathcal{N}=4$. Now the Coulomb limit is simply obtained by taking $T=1$ at the $A$-twisted point (or $T\to\mathfrak{q}^{1/2}$ at the untwisted point). For the Higgs limit, we should take $\lim_{\mathfrak{q}\to 0}\mathcal{I}_{SCI}(\mathfrak{q},T)|_{T\rightarrow\mathfrak{q}^{-\frac{1}{2}}T}$. This gives the Coulomb and Higgs limits of the index to be
@@
which is consistent with the expectation that both the Coulomb and Higgs branches are trivial. This strongly supports our identification of $U(1)_{A}$.
Now, we perform F-maximization to show that the superconformal $R$ symmetry is at $\nu=0$. We compute the three-sphere partition function using the Bethe ansatz equation (BAE) method of Closset:2017 instead of the ordinary integral formula, since it offers computational advantages. From the 3d gauge theory, we construct the effective twisted superpotential $W$ as a function of the mixing parameter $\nu$. The Bethe ansatz equation $\{\exp(2\pi i\partial W(u)/\partial u_{i})=1\}$ for this theory turns out to have two solutions $u_{0}$ and $u_{1}$. The three sphere partition function is now given by
$Z_{S^{3}}=\sum_{u_{0},u_{1}}\mathcal{H}^{-1}(u)\mathcal{F}^{1}(u).$ (63)
We plot the $S^{3}$ partition function as a function of mixing parameter $\nu$ in Figure 3. We see that the absolute value $|Z_{S^{3}}|$ is minimized (or $F$ is maximized) at $\nu=0$ so that it indeed corresponds to the superconformal point, which further corroborates our identification of $U(1)_{A}$.
Figure 3: The absolute value of the three sphere partition function of the 3d theory for $(A_{1},A_{2})$ as a function of mixing parameter $\nu$. It is minimized at $\nu=0$.
Now, armed with $U(1)_{A}$ symmetry, we can compute the $T$-refined half index. We have
@@
where we see that the expression in the second line agrees with our proposal (28). This expression can be shown to agree with the fermionic sum formula for the Macdonald index of (18) for this case Gaiotto:2024ioj .

### v2

@@ 3 Macdonald index from the twisted compactification
    Extract the (matter contents and their charges of) 3d $\mathcal{N}=2$ Abelian Chern-Simons matter theory from the above expression for the half-index.
@@
    In general, this gauge theory admits a superpotential deformation, which triggers an RG flow to $\mathcal{N}=4$ SCFT in the IR. The superpotential is expected to break most of the global symmetry except for the distinguished $U(1)_{A}$ symmetry (and Higgs branch flavor symmetries of 4d), which becomes part of the $SO(4)_{R}$ symmetry in the IR. The task is to identify this $U(1)_{A}$ symmetry.
@@
    Refine the half-index by turning on the fugacity $T$ for $U(1)_{A}$. This gives the Macdonald index of the 4d theory.
@@
The starting point of our procedure is an observation that the fermionic sum formula from the IR formula can be viewed as a twisted half-index of a 3d $\mathcal{N}=2$ Abelian Chern-Simons matter theory (ACSM) Gaiotto:2024ioj . The half-index of a 3d $\mathcal{N}=2$ theory with a supersymmetric boundary condition $B$ is defined by Gadde:2013wq ; Gadde:2013sca ; Yoshida:2014ssa ; Dimofte:2017
@@
Consider a $U(1)^{N}$ Chern-Simons theory with level matrix $K$ and $M$ chiral multiplets $\Phi_{\alpha}$ with $U(1)_{i}$ gauge charge $Q_{i\alpha}$, and $R$-charge $R_{\alpha}$. Imposing the supersymmetric Dirichlet boundary conditions to all the vector and chiral multiplets, we obtain the half-index Dimofte:2017
$I\!\!I\!(q,{\boldsymbol{x}})=\frac{1}{(q;q)_{\infty}^{N}}\sum_{\mathbf{m}\in\mathbb{Z}^{N}}q^{\frac{1}{2}\mathbf{m}^{t}K\mathbf{m}+\frac{1}{2}\boldsymbol{\rho}\cdot\mathbf{m}}{\boldsymbol{x}}^{K{\bf m}}\prod_{\alpha=1}^{M}\left(q^{R_{\alpha}/2}q^{1-{\mathbf{m}}\cdot\boldsymbol{Q}_{\alpha}}{\boldsymbol{x}}^{-\boldsymbol{Q}_{\alpha}};q\right)_{\infty}\ ,$ (39)
where $\rho_{i}$ denotes the mixed Chern-Simons level between the gauge group $U(1)_{i}$ and background $U(1)_{R}$ symmetry, and $x_{i}$’s are fugacity for the boundary current $U(1)^{N}_{\partial}$. Here we use the standard abbreviation for the multiple fugacities for the symmetry as $\boldsymbol{x}^{\boldsymbol{m}}\equiv\prod_{i}x_{i}^{m_{i}}$. In most of the examples below, we consider the fully unrefined index with ${\boldsymbol{x}}\rightarrow 1$.
Upon comparing this expression with the one obtained from the IR trace formula for the Schur index (23), we extract an abelian Chern-Simons theory. For example, the Schur index of the form (26) for the $(A_{1},A_{2N})$ theory gives rise to $U(1)^{2N}$ Chern-Simons (CS) theory, where each $U(1)$ is coupled to 2 chiral multiplets of charge $1$, giving a total of $M=4N$ chiral multiplets. We can also read the CS level matrix $K$, mixed U(1)RU(1)-R CS level $\rho$ for each $U(1)$ gauge factors, and $R$-charges for the chiral multiplets by comparing the formulae.
This description of the abelian Chern-Simons theory can be motivated by considering a Janus configuration in the Coulomb branch of the 4d SCFT, induced by the $U(1)_{r}$ twisting. See section 6 of Gaiotto:2024ioj for further discussion.
@@
The abelian Chern-Simons theory obtained in this way is expected to flow to an SCFT with $\mathcal{N}=4$ supersymmetry enhancement, after a suitable superpotential deformation. The IR $\mathcal{N}=4$ SCFT has an $\mathcal{N}=2$ superconformal subalgebra where $U(1)_{R}$ symmetry embeds into the full $\mathcal{N}=4$ R-symmetry group $SU(2)_{H}\times SU(2)_{C}$, leaving $U(1)_{A}=U(1)_{H}-U(1)_{C}$ as a distinguished global symmetry. When the 3d $\mathcal{N}=4$ SCFT is expected to have a zero-dimensional Coulomb/Higgs branch, we anticipate that there exists a suitable superpotential deformation that lifts most of the global symmetry of the UV ACSM. This leaves one combination of $U(1)$, which can be identified with the IR $U(1)_{A}$ symmetry. To find the superconformal $R$-symmetry, we consider a one-parameter mixing of the $R$-symmetry
@@
and perform $F$-maximization Jafferis:2010un . We define our normalization so that $\nu=0$ corresponds to the superconformal $R$-symmetry $U(1)_{R_{0}}=U(1)_{R}=U(1)_{C}+U(1)_{H}$.
@@
In the special limit $\nu=-1$, the half-index computes the index with the twisted spin $J+R_{\nu}/2=J+R_{H}$, realizing the topological $A$-twist. We therefore define
@@
The 3d $\mathcal{N}=4$ SCFTs obtained by twisted compactification of $(A_{1},G)$ Argyres-Douglas theories are expected to have zero-dimensional Coulomb branches, while their Higgs branches coincide with those of the original 4d SCFT. On the other hand, the proposed $\mathcal{N}=2$ ACSM descriptions generally exhibit a larger global symmetry than the Higgs branch symmetry. We therefore expect that there exists an appropriate superpotential deformation that breaks the global symmetries to $U(1)_{A}\times T_{H}$, where $T_{H}\subset G_{H}$ denotes the Cartan subgroup of the Higgs branch symmetry.
@@
    The superconformal index contains a term $q^{3/2}(T+T^{-1})$, signaling the existence of the extra supercurrent multiplets in the $\mathcal{N}=4$ algebra [^2] This condition is neither necessary nor sufficient, since there may exist, for example, chiral primary operators with $R$-charge 3. In fact, we have examples where supersymmetry enhancement is expected, yet this term is partially or entirely absent., as well as a possible moment map operator for the Higgs branch symmetry appearing at $q^{1/2}$ order (when such a symmetry is present).
@@
    The Coulomb branch Hilbert series is trivial. [^3] For a 3d SCFT, the superconformal index can be written as $\mathcal{I}(\mathfrak{q},\mathfrak{t})=\mathrm{tr}_{\mathbb{S}^{2}}(-1)^{F}\mathfrak{q}^{j_{3}+\frac{1}{2}(R_{H}+R_{c})}\mathfrak{t}^{R_{H}-R_{C}}$ (44) and it gets contribution only from states satisfying $E-R_{H}-R_{C}-j_{3}=0$. We can rewrite the index by $\mathcal{I}(\mathfrak{q},\mathfrak{t})=\mathrm{tr}(-1)^{F}(\mathfrak{q}^{\frac{1}{2}}\mathfrak{t})^{E-R_{c}}(\mathfrak{q}^{\frac{1}{2}}\mathfrak{t}^{-1})^{E-R_{H}}.$ (45) The Higgs and Coulomb branch Hilbert series are obtained by taking the limit Razamat:2014 Coulomb: $\qquad\mathfrak{q}\rightarrow 0,\mathfrak{t}\rightarrow 0,\qquad\mathfrak{q}^{\frac{1}{2}}\mathfrak{t}^{-1}\;\text{fixed}$ (46a) Higgs: $\qquad\mathfrak{q}\rightarrow 0,\mathfrak{t}\rightarrow\infty,\qquad\mathfrak{q}^{\frac{1}{2}}\mathfrak{t}\;\text{fixed}$ (46b) which give extra shortening condition $E=R_{H}$ and $E=R_{C}$ for the trace, respectively.
@@
    The Higgs branch Hilbert series agrees with that of the Higgs branch chiral ring of 4d SCFT.
In some low-rank examples, this procedure determines the $U(1)_{A}$ symmetry in the ACSM description, and one can check that the refined half-index (42) reproduces the conjectural Macdonald index formula described in section 2.3.
One subtlety in identifying $U(1)_{A}$ arises from the fact that a global symmetry can be redefined by mixing it with gauge symmetries. The Dirichlet half-index is sensitive to this choice because, at the boundary, the gauge symmetry is reduced to a global symmetry. Therefore, the computation of the refined half-index is not completely canonical and depends on this choice. In the examples below, we show that there exists a representation of the $U(1)_{A}$ symmetry that reproduces the conjectural Macdonald index presented in section 2.3, providing a strong consistency check despite its non-uniqueness.
For higher rank examples, where F-maximization is not practically feasible, we begin with the assumption that the Schur index coincides with the A-twisted half-index of the 3d $\mathcal{N}=4$ SCFT, and that the conjectural Macdonald index introduced in section 2.3 computes its refined half-index. This assumption completely determines the data for the superconformal R-symmetry. We then compute the corresponding superconformal index and the Coulomb and Higgs branch Hilbert series, and test them against the conditions described above.
@@ 4 Examples
In this section, we apply our procedure for $(A_{1},G)$ type Argyres-Douglas theories to recover our conjectural closed-form expression for the Macdonald index.
@@ 4.1 $(A_{1},A_{2n})$ theory
The BPS quiver for the $(A_{1},A_{2n})$ theory is given in Figure 2.
@@
This describes the canonical chamber with an equal number of sources and sinks. The trace formula (23) gives the Schur index Cordova:2015nma ,
@@
where $b^{g}_{ij}=2\delta_{ij}-C[g]_{ij}$ and $C[g]$ is the Cartan matrix of $g$. This expression agrees with the character of the Virasoro minimal model $M(2,2n+3)$ Cecotti:2015lab .
We begin with the observation that this formula coincides with the Dirichlet half-index of 3d $U(1)^{2n}$ Chern-Simons theory with level
$K=b^{A_{2n}}=\begin{pmatrix}0&1&0&0&\cdots&0&0\\ 1&0&0&0&\cdots&0&0\\ 0&0&0&1&\cdots&0&0\\ 0&0&1&0&\cdots&0&0\\ &&&\vdots\\ 0&0&0&0&\cdots&0&1\\ 0&0&0&0&\cdots&1&0\end{pmatrix}\$. (48)
Additionally, we have two chiral multiplets $\Phi_{i}$ and $\widetilde{\Phi}_{i}$ with charge 1-1 for each $U(1)_{i}\subset U(1)^{2n}$, giving a total of $4n$ chiral multiplets. Therefore, we can write the matter content and the charge matrix $Q_{i\alpha}$ as
@@
The half-index formula shows that the $R$-charges for all chiral multiplets are zero and the mixed Chern-Simons level between $U(1)_{R}$ and $U(1)_{i}$ is $\rho_{i}=2$ for all $i=1,2,\cdots,2n$. Before adding the superpotential term, the ACSM theory has $U(1)^{4n}$ flavor symmetry.
We expect that there exists a superpotential deformation that breaks all the symmetries except for one $U(1)$, which we denote by $U(1)_{A}$. Our task is to identify this symmetry. We further expect that the Schur index (47), when interpreted as the half-index of the 3d $\mathcal{N}=4$ theory, is computed in the A-twisted limit.
@@ 4.1.1 $A_{2}$
Let us start with the simplest case of $n=1$. The trace formula gives
@@
This expression can be identified with the half-index of the 3d $\mathcal{N}=2$ Abelian Chern-Simons theory with $U(1)^{2}$ gauge group, with the following data:
@@
This gauge theory has $U(1)^{4}$ flavor symmetry, corresponding to the rotations of each chiral multiplet. We expect that there exists an appropriate superpotential deformation constructed from dressed monopole operators which break $U(1)^{4}$ flavor symmetry down to the single $U(1)_{A}$. Such monopole operators should have $R$-charge 2, and there are four of them in total, as listed in Table 1.
@@
Consider the following superpotential [^4] Any choice of three out of four monopole operators gives rises the same result, thanks to the symmetry.
@@
The resulting unbroken global symmetry is $U(1)_{\widetilde{\Phi}_{1}}$, which rotates $\widetilde{\Phi}_{1}$ with charge +1+1 and keeps the rest intact. By mixing with gauge symmetry, this global symmetry can also be represented as a combination of the topological symmetry for the second gauge node and the symmetry that rotates $\Phi_{1}$ with charge 1-1.555Consider the bosonic action $2A_{1}dA_{2}+|D_{(-A_{1}+A^{\text{bg}})}\widetilde{\Phi}|^{2}+|D_{(-A_{1})}\Phi|^{2}+\cdots$. By redefining the gauge connection $A_{1}\rightarrow A_{1}+A^{\text{bg}}$, we obtain $2A_{1}dA_{2}+2A^{(\text{bg})}dA_{2}+|D_{(-A_{1})}\widetilde{\Phi}|^{2}+|D_{(-A_{1}-A^{(bg)})}\Phi|^{2}+\cdots$ We will now argue that this is the desired superpotential and verify that this unbroken symmetry is the axial symmetry $U(1)_{A}$ that accounts for the Macdonald grading.
We can now compute the superconformal index of the 3d SCFT via the gauge theory description, refined to include the candidate $U(1)_{A}$ fugacity. In order to determine the superconformal R-charge, we perform F-maximization,via the three-sphere partition function using the Bethe ansatz method of Closset:2017 . From the 3d gauge theory, we construct the effective twisted superpotential $W$ as a function of the mixing parameter $\nu$. The Bethe ansatz equation $\{\exp(2\pi i\partial W(u)/\partial u_{i})=1\}$ for this theory turns out to have two solutions $u_{0}$ and $u_{1}$. The three-sphere partition function is then given by
$Z_{S^{3}}=\sum_{u_{0},u_{1}}\mathcal{H}^{-1}(u)\mathcal{F}^{1}(u).$ (60)
We plot the $S^{3}$ partition function as a function of mixing parameter $\nu$ in Figure 3. We see that the absolute value $|Z_{S^{3}}|$ is minimized (or $F$ is maximized) at $\nu=0$, which corresponds to the superconformal point.
Figure 3: The absolute value of the three-sphere partition function of the 3d theory for $(A_{1},A_{2})$ as a function of mixing parameter $\nu$. The function is minimized at $\nu=0$.
The superconformal index at the fixed point is
@@
The presence of the term $(T+T^{-1})\mathfrak{q}^{\frac{3}{2}}$ provides a strong evidence for the $\mathcal{N}=4$ supersymmetry enhancement. Now the Coulomb limit is simply obtained by taking the limit $T\to\mathfrak{q}^{1/2}$. For the Higgs limit, we take the limit $\lim_{\mathfrak{q}\to 0}\mathcal{I}_{SCI}(\mathfrak{q},T)|_{T\rightarrow\mathfrak{q}^{-\frac{1}{2}}T}$. We find
@@
which is consistent with the expectation that both the Coulomb and Higgs branches of the 3d SCFT are trivial. This strongly supports our identification of $U(1)_{A}$.
Now, armed with $U(1)_{A}$ symmetry, we can compute the $T$-refined half-index, in the A-twisted limit. This gives
@@
where we see that the expression in the second line reproduces with our proposal (30). This expression can be shown to agree with the fermionic sum formula (18) for the Macdonald index for this case Gaiotto:2024ioj .